# A detailed study on ergodic to non-ergodic phase transition in the dissipative anisotropic Dicke model

### v1

@@ Acknowledgments
We sincerely thank Auditya Sharma and Devendra Singh Bhakuni for fruitful discussions and constructive comments. We are also grateful to the High-Performance Computing (HPC) facility at IISER Bhopal, where large-scale calculations in this project were run. P.D. acknowledges IISERB for the PhD fellowship. This research was supported in part by the International Centre for Theoretical Sciences (ICTS) for participating in the program Stability of Quantum Matter in and out of Equilibrium at Various Scales (ICTS/SQMVS2024/01).
@@ Appendix A Eigenvalue properties
We give a review of the distribution function of the eigenvalues and levels spacings ratio of a many-body quantum system, which motivates us to study the nature of such function for this model. Based on the eigenvalue distribution of a Hermitian system, it can be broadly classified into two categories: (i) regular/integrable/non-ergodic and (ii) chaotic/ergodic. Berry and Tabor hypothesized Berry and Tabor (1977) that, for integrable systems, the levels are uncorrelated, and hence the level spacing distribution follows the one-dimensional Poisson distribution. However for correlated spectrum, the famous Wigner ‘surmise’ Bohr and Mottelson (1969) is a powerful tool to investigate the spacing distribution. In $1984$ Bohigas, Giannoni, and Schmit (BGS) hypothesized Bohigas et al. (1984) that chaotic systems would follow statistical behavior described by random matrix theory appropriate to their symmetry classes. Following BGS, Grobe, Haake, and Sommers (GHS) Grobe et al. (1988) expanded the above spectral contrast between integrability and chaos for Markovian open quantum systems. They found that in the integrable limit, the nearest neighbor spacing distribution of radial distance $s$ for the complex eigenvalues of Liouvillian agree with the two-dimensional (2D) Poisson distribution. Such distribution is the same as Wigner’s conjecture for spacing distribution in the Gaussian orthogonal ensemble (GOE). On the other hand for a fully chaotic limit, the spacing distribution agrees with the Ginibre ensemble  Ginibre (1965); Grobe et al. (1988); Markum et al. (1999); Haake (1991); Hamazaki et al. (2020).
Here we study the eigenvalue properties of the Liouvillian $\mathcal{L}$. Since complex eigenvalues are involved, we consider generalized notions of the level spacing distribution and the consecutive level spacing ratio  Prasad et al. (2022) . By analyzing these quantities as a function of the coupling parameters, we try to sketch the phase diagram which carries information on the non-ergodic to ergodic phase transition in the dissipative ADM.
@@
The eigenvalues of $\mathcal{L}$ lie in the nonpositive plane, which prevents the divergence of the Eq.(3). The non-real eigenvalues possess complex conjugates, demonstrating the hermiticity of the density matrix. The zero eigenvalue of $\mathcal{L}$ corresponds to a steady state. The real part of the complex eigenvalues corresponds to the loss or decay of the initial density matrix Minganti et al. (2018).
In Fig. 6 we show the scatter plot of the complex spectrum of the Liouvillian $\mathcal{L}$ for two different set of $(g_{1},g_{2})$ pairs: $g_{1}=1.25$ and (a) $g_{2}=0.1$, (b) $g_{2}=1.0$ corresponding to the non-ergodic and ergodic regimes respectively. They clearly show the different natures of the spectrum in the regular and chaotic regimes respectively. For a clearer understanding of the two different phases, one should study the level spacing statistics. Additionally, we would like to discuss briefly how to calculate the level spacing distribution for a non-Hermitian system.
We first carry out an unfolding of the spectrum to remove the system-specific aspects of the level spacing statistics. The complex level spacing distribution $P(s)$ is then produced by using the unfolded spectrum to create a histogram of the Euclidean distance $s$ between nearest-neighbor eigenvalues in the complex plane. For the unfolding process we mainly follow the references Haake (1991); Akemann et al. (2019); Hamazaki et al. (2020).
@@
Figure 7: Level-spacing distribution of the complex spectrum of the Liouvillian $\mathcal{L}$ in (a) the non-ergodic phase with $g_{1}=1.25,\ g_{2}=0.1$ and (b) the ergodic phase with $g_{1}=1.25,\ g_{2}=1.0$. We find remarkable agreement with the 2D Poisson distribution $P_{\text{2D-P}}(s)$ in Eq. and that of the GinUE RMT prediction $P_{\text{GinUE}}(s)$. The atom number $N=10$, bosonic cut-off: $n_{\text{max}}=26$.
@@
On the other hand, for the fully chaotic (or ergodic) limit the spacing distribution agrees with the corresponding distribution of the Ginibre ensemble of complex Gaussian non-Hermitian random matrices, given by:
@@
Figure 8: Scatter plot of the complex level-spacing ratio $z$ introduced in Eq. 18 for the atom number $N=10$, for (a) a point in the non-ergodic phase: $g_{1}=1.25,\ g_{2}=0.1$, which shows the isotropic behavior and (b) a point in the ergodic pgase: $g_{1}=1.25,\ g_{2}=1.0$, which is not isotropic. The bosonic cut-off: $n_{\text{max}}=26$.
[…]
We study the level spacing statistics separately for these two pair points in different phases (see Fig. 7). Fig. 7 displays the striking agreement between the distributions calculated from the spectrum of $\mathcal{L}$ and the $2D$ Poisson distribution in the integrable or regular or the non-ergodic regime and the GinUE forecast in the chaotic or ergodic regime, respectively. In the ergodic phase $P(s)\approx s^{3}$, which indicates the chaotic behavior of the complex eigenvalues as they repel each other and this is consistent with Ginibre distribution in Eq. 16. On the other hand in the non-ergodic phase, we have $P(s)\approx s$ indicating the integrable nature of this regime and consistent with $2D$ Poisson distribution in the complex plane.
@@ A.1 Consecutive level spacing ratio
The complex level spacing ratio Sá et al. (2020); Prasad et al. (2022) can be defined as:
@@
where the superscripts NN (NNN) stand for nearest (next nearest) neighbor. It captures next-nearest neighbor information that is overlooked in the levelspacing statistics. The advantage of the calculation is we don’t require the unfolding procedure. In Fig. 8 we show the scatter plot of $z$ in the non-ergodic ($g_{1}=1.25,\ g_{2}=0.1$) and the ergodic phases ($g_{1}=1.25,\ g_{2}=1.0$) separately. For the non-ergodic case, the scatter plot is symmetric and isotropic whereas for the ergodic case, it is anisotropic which is another signature for the connection to RMT Sá et al. (2020). In Fig. 9 we show the absolute value of the average level spacing ratio for this model. The figure shows a non-ergodic phase to ergodic phase transition similar to the closed ADM Das et al. (2023a). Here in the non-ergodic phase, $\langle r\rangle\approx 0.69$, (dark color) which is equivalent to the system following the 2D Poisson distribution, whereas in the ergodic phase, $\langle r\rangle\approx 0.74$ (light color), and it corresponds to the Ginibre unitary ensemble.

### v2

@@ Acknowledgments
We sincerely thank Auditya Sharma and Devendra Singh Bhakuni for fruitful discussions and constructive comments. We are also grateful to the High-Performance Computing (HPC) facility at IISER Bhopal. P. D. acknowledges IISERB for providing the PhD fellowship that supported the initial phase of this project. P. D. also acknowledges support from the European Research Council (ERC) through a postdoctoral fellowship (Project No. N1-0273). S. S. thanks the University Grants Commission for providing a PhD fellowship (Student ID: MAY 2018–528071), which supported the initial phase of this research. SS also acknowledges research support and funding from BERLIN QUANTUM, an initiative endowed by the Innovation Promotion Fund of the city of Berlin, and the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) via the Research Unit FOR 5688 (Project No. 521530974). We also thank the International Centre for Theoretical Sciences (ICTS) for hosting the program Stability of Quantum Matter in and out of Equilibrium at Various Scales (ICTS/SQMVS2024/01), where this work and the associated discussions were initiated.
@@ Appendix A Eigenvalue properties
We provide a brief review of the distribution function of the eigenvalues and level spacing ratios of a many-body quantum system, which motivates us to study the nature of such a function for this model. Based on the eigenvalue distribution of a Hermitian system, it can be broadly classified into two categories: (i) regular/integrable/non-ergodic and (ii) chaotic/ergodic. Berry and Tabor hypothesized Berry and Tabor (1977) that, for integrable systems, the levels are uncorrelated, and hence the level spacing distribution follows the one-dimensional Poisson distribution. However, for correlated spectrum, the famous Wigner ‘surmise’ Bohr and Mottelson (1969) is a powerful tool to investigate the spacing distribution. In $1984$ Bohigas, Giannoni, and Schmit (BGS) hypothesized Bohigas et al. (1984) that chaotic systems would follow statistical behavior described by random matrix theory appropriate to their symmetry classes. Following BGS, Grobe, Haake, and Sommers (GHS) Grobe et al. (1988) expanded the above spectral contrast between integrability and chaos for Markovian open quantum systems. They found that in the integrable limit, the nearest neighbor spacing distribution of radial distance $s$ for the complex eigenvalues of Liouvillian agrees with the two-dimensional (2D) Poisson distribution. Such a distribution is the same as Wigner’s conjecture for the spacing distribution in the Gaussian orthogonal ensemble (GOE). On the other hand, for a fully chaotic limit, the spacing distribution agrees with the Ginibre ensemble  Ginibre (1965); Grobe et al. (1988); Markum et al. (1999); Haake (1991); Hamazaki et al. (2020).
Here, we study the eigenvalue properties of the Liouvillian $\mathcal{L}$. Since complex eigenvalues are involved, we consider generalized notions of the level spacing distribution and the consecutive level spacing ratio  Prasad et al. (2022) . By analyzing these quantities as a function of the coupling parameters, we try to sketch the phase diagram, which carries information on the non-ergodic to ergodic phase transition in the dissipative ADM.
@@
The eigenvalues of $\mathcal{L}$ lie in the nonpositive plane, which prevents the divergence of Eq.(3). The non-real eigenvalues possess complex conjugates, demonstrating the hermiticity of the density matrix. The zero eigenvalue of $\mathcal{L}$ corresponds to a steady state. The real part of the complex eigenvalues corresponds to the loss or decay of the initial density matrix Minganti et al. (2018).
In Fig. 4, we show the scatter plot of the complex spectrum of the Liouvillian $\mathcal{L}$ for two different set of $(g_{1},g_{2})$ pairs: $g_{1}=1.25$ and (a) $g_{2}=0.1$, (b) $g_{2}=1.0$ corresponding to the non-ergodic and ergodic regimes respectively. They clearly show the different natures of the spectrum in the regular and chaotic regimes, respectively. For a clearer understanding of the two different phases, one should study the level spacing statistics. Additionally, we would like to discuss briefly how to calculate the level spacing distribution for a non-Hermitian system.
We first carry out an unfolding of the spectrum to remove the system-specific aspects of the level spacing statistics. The complex level spacing distribution $P(s)$ is then produced by using the unfolded spectrum to create a histogram of the Euclidean distance $s$ between nearest-neighbor eigenvalues in the complex plane. For the unfolding process, we mainly follow the references Haake (1991); Akemann et al. (2019); Hamazaki et al. (2020).
@@
Figure 5: Level-spacing distribution of the complex spectrum of the Liouvillian $\mathcal{L}$ in (a) the non-ergodic phase with $g_{1}=1.25,\ g_{2}=0.1$ and (b) the ergodic phase with $g_{1}=1.25,\ g_{2}=1.0$. We find remarkable agreement with the 2D Poisson distribution $P_{\text{2D-P}}(s)$ in Eq. and that of the GinUE RMT prediction $P_{\text{GinUE}}(s)$. The atom number $N=10$, bosonic cut-off: $n_{\text{max}}=26$. The other parameters are: $\omega=\omega_{0}=1$, $\kappa=1$.
@@
On the other hand, for the fully chaotic (or ergodic) limit, the spacing distribution agrees with the corresponding distribution of the Ginibre ensemble of complex Gaussian non-Hermitian random matrices, given by:
@@
Figure 6: Scatter plot of the complex level-spacing ratio $z$ introduced in Eq. 16 for the atom number $N=10$, for (a) a point in the non-ergodic phase: $g_{1}=1.25,\ g_{2}=0.1$, which shows the isotropic behavior and (b) a point in the ergodic pgase: $g_{1}=1.25,\ g_{2}=1.0$, which is not isotropic. The bosonic cut-off: $n_{\text{max}}=26$. The other parameters are: $\omega=\omega_{0}=1$, $\kappa=1$.
and $\bar{s}=\int_{0}^{\infty}s\bar{P}_{\text{GinUE}}(s)\hskip 2.84526ptds$, $\Gamma(1+k,s^{2})=\int_{s^{2}}^{\infty}t^{k}e^{-t}dt$, is the incomplete Gamma function.
We study the level spacing statistics separately for these two pair points in different phases (see Fig. 5). Fig. 5 displays the striking agreement between the distributions calculated from the spectrum of $\mathcal{L}$ and the $2D$ Poisson distribution in the integrable or regular or the non-ergodic regime and the GinUE forecast in the chaotic or ergodic regime, respectively. In the ergodic phase $P(s)\approx s^{3}$, which indicates the chaotic behavior of the complex eigenvalues as they repel each other, and this is consistent with Ginibre distribution in Eq. 14. On the other hand, in the non-ergodic phase, we have $P(s)\approx s$ indicating the integrable nature of this regime and consistent with $2D$ Poisson distribution in the complex plane.
@@ A.1 Consecutive level spacing ratio
The complex level spacing ratio et al. (2020); Prasad et al. (2022) can be defined as:
@@
where the superscripts NN (NNN) stand for the nearest (next nearest) neighbor. It captures the information about the next closest neighbor that is overlooked in the level spacing statistics. The advantage of the calculation is that we don’t require the unfolding procedure. In Fig. 6, we show the scatter plot of $z$ in the non-ergodic ($g_{1}=1.25,\ g_{2}=0.1$) and ergodic phases ($g_{1}=1.25,\ g_{2}=1.0$) separately. For the non-ergodic case, the scatter plot is symmetric and isotropic, whereas for the ergodic case, it is anisotropic, which is another signature for the connection to RMT et al. (2020). In Fig. 7 we show the absolute value of the average level spacing ratio for this model. The figure shows a non-ergodic phase to ergodic phase transition similar to the closed ADM Das et al. (2023a). Here in the non-ergodic phase, $\langle r\rangle\approx 0.69$, (dark color) which is equivalent to the system following the 2D Poisson distribution, whereas in the ergodic phase, $\langle r\rangle\approx 0.74$ (light color), and it corresponds to the Ginibre unitary ensemble.
Appendix B Dependence of the exponent $z$ by varying $g_{1}$ and $g_{2}$.
We provide a contour plot of the exponent $z$ by varying $g_{1}$ and $g_{2}$ for $\kappa=0.1$.
Figure 8: The exponent $z$ (extracted from the scaling of Liouvillian gap $\Delta$ as a function of atom number) as a function of $g_{1}$ and $g_{2}$ for $\kappa=0.1$ and $n_{\text{max}}=26$. Here, the atom number NN is changing from $4:10$. For varying $g_{1}$ and $g_{2}$, the system shows a response which has similarities to ENET. For example, in the vicinity of $g_{1}=0$ and $g_{2}=0$ (non-ergodic phase), the value of $z$ (dark-orange portion) is different from the remaining part (ergodic phase).
Our results successfully capture the key features of the system, even though the extreme point ($g_{1}=g_{2}$), which occurs farthest from the integrability, is not yet clearly visible. Nevertheless, the dependence of $z$ on $g_{1}$ and $g_{2}$ shows qualitative similarities to ENET. To achieve a clearer understanding, simulations with a larger number of atoms and a higher bosonic cutoff will be required.
Appendix C Average participation ratio for different $\kappa$
Here we provide the contour plots of $P^{\rm LR}_{\rm avg}$ for different $\kappa$ by varying $g_{1}$ and $g_{2}$ in Fig. 9. We observe that the study of $P^{\rm LR}_{\rm avg}$ successfully captures the ENET, and for changing $\kappa$, the qualitative features remain unchanged.
Figure 9: Average of participation ratio ($P^{\rm LR}_{\text{avg}}$), using biorthonormal basis considering both left and right eigenvectors of the Liouvillian for dissipative ADM as a function of the rotating and counter-rotating coupling parameters $g_{1}$ and $g_{2}$. The atom number: $N=4$, the bosonic cut-off: $n_{\text{max}}=10$. For this figures $\omega_{0}=\omega=1,\ \kappa=1.0,\,0.1,\,0.01$ respectively.